A race car starts from rest and accelerates at a constant rate and reaches a speed of 160kmh-¹ in 6.0seconds. It continues at this speed for another 5 seconds. What is the car’s average speed during the first 11 seconds?
1
Expert's answer
2019-06-07T11:45:52-0400
"d_1=0.5vt_1=0.5\frac{160}{3.6}(6)=\frac{400}{3} m"

"d_2=vt_2=\frac{160}{3.6}(5)=\frac{2000}{9} m"

"V=\frac{d_1+d_2}{t_1+t_2}=\frac{\frac{400}{3}+\frac{2000}{9}}{6+5}=32\frac{m}{s}"